OpenStudy (anonymous):

There are 20 parrots at the animal sanctuary. Their population is increasing at a rate of 15% per year. There are also 24 snakes at the sanctuary. Each year 4 more snakes are born. Part A: Write functions to represent the numbers of parrots and snakes at the sanctuary throughout the years. Part B: How many parrots are at the sanctuary after 10 years? How many snakes are at the sanctuary after the same number of years? Assume there are no deaths to the animals during this time. Part C: After approximately how many years is the number of parrots and snakes the same?

zepdrix (zepdrix):

Hey there :) So after 1 year, the parrot population will be the base amount, 20 parrots, plus 15% of that 20. 1 year = 20 + 15% * 20 Oh oh oh, but then the next year, our population increase should be based on the new population amount. Hmm this is going to be tricky. I thought it was a simple linear equation, but it looks like no. 2 years = 20 + 15%(1 year amount) 2 years = 20 + 15%(20 + 15% * 20)

zepdrix (zepdrix):

Have you learned about exponential functions yet? I should ask because maybe I'm misreading the question :)

OpenStudy (anonymous):

I was reading the lesson but got confused.It's really hard when you can't contact a teacher that understands I'm not a fast learner.

OpenStudy (anonymous):

i calculated what 20+15%=23+20=460

OpenStudy (mathmale):

Zepdrix has asked you an important question: Have you studied "exponential functions" yet?

OpenStudy (mathmale):

Thank you for sharing your work: "i calculated what 20+15%=23+20=460" .... but I need to ask you to explain this. Having more than one = symbol in one equation can be a sign of trouble.

OpenStudy (anonymous):

ill do it by myself.

OpenStudy (mathmale):

Do y ou have a textbook or workbook for this course? If so, what is the title of the chapter you're studying?